SPECTROSCOPY OF LARGE VAN DER WAALS CLUSTERS IN HELIUM DROPLETS

SUSUMU KUMA, HARUKA GOTO, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan; \underline{TAKAMASA MOMOSE}, Department of Chemistry, University of British Columbia, Vancouver, BC V6T 1Z1, Canada; and ANDREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90080, USA.

Large clusters of Ar_{N} and (H_{2})_{N} (N = 0-1500) have been prepared in He droplets. The clusters have been studied via the S_{1} ← S_{0} LIF spectra of the attached phthalocyanine (Pc) molecules. The band origin of Pc shifts towards low frequency upon increase of the cluster size. The amount of the shift was found to be Δν̃ = −400cm^{−1} in the case when Ar atoms were captured by He droplets after pick up of Pc molecules. Two times smaller shift was obtained with the reversed order of the pick up events, indicating Pc molecules reside on the surface of the Ar clusters. On the other hand, the same shift was obtained in pH_{2} clusters irrespective of the order of pick up, which shows Pc molecules penetrate inside the pH_{2} clusters at low temperature of He droplets of 0.38 K.